A random variable of size 30 bin(20,0.6) ie binomial distribution.Find (x<12.2) (x>12.2(


1
Expert's answer
2021-12-14T01:17:48-0500

Given:

"\begin{aligned}

&X \sim \operatorname{Binomial}(20,0.6) \\

&N=20 \\

&p=0.06

\end{aligned}"

Checking for normality condition

"\begin{aligned}

&N p=20 \times 0.6=12 \\

&N(1-p)=20 \times(1-0.6)=8

\end{aligned}"

Since both "\mathrm{Np}" and "\mathrm{N}(1-\mathrm{p})" is greater than 5, the given binomial distribution can be approximated to the normal distribution.

The normal approximation for the sample of size "\mathrm{n}=30" is given as

"\begin{aligned}

\bar{X} & \sim N\left(\mu, \frac{\sigma}{\sqrt{n}}\right) \\

\mu_{X} &=N p \\

&=20 \times 0.6 \\

&=12 \\

\sigma_{X} &=\frac{\sqrt{N p(1-p)}}{\sqrt{n}} \\

&=\frac{\sqrt{20(0.6)(1-0.6)}}{\sqrt{30}} \\

&=0.4

\end{aligned}"

i)

"\begin{aligned}
& \text{The required probability is calculated as}\\
& P(\bar{X}<12.2) =P\left(\frac{\bar{X}-\mu}{\sigma}-\frac{12.2-\mu}{\sigma}\right) \\

&=P\left(Z<\frac{12.2-12}{0.4}\right) \\

&=P(Z<0.5) \\

&=0.69146(\text { Using the standard normal table })\\
& \text{Thus, the required probability is 0.69146.}
\end{aligned}"

ii)

The required probability is calculated as

"\begin{aligned}

P(\bar{X}>12.2) &=1-P(\bar{X}<12.2) \\

&=1-P\left(\frac{\bar{X}-\mu}{\sigma}-\frac{12.2-\mu}{\sigma}\right) \\

&=1-P\left(Z<\frac{12.2-12}{0.4}\right) \\

&=1-P(Z<\frac{1}{2}) \\

&=1-0.69146(\text { Using the standard normal table }) \\

&=0.30854

\end{aligned}"

Thus, the required probability is "0.30854 ."